MATH 135 – Calculus 1 Derivative Rule Practice October 11, 2013

Directions: Use the sum, constant multiple, product, and quotient rules to compute each of the following derivatives. Say which rules you are using for each one.

I. Find f′(x) for

f(x) = x^{4}e^{x} − \frac{3}{\sqrt{x}}

II. Find G′(t) for

G(t) = e^{−t}(t^{2} − 4t + 3)

III. Find H′(z) and H′′(z) for

H(z) = \frac{z^{2}}{z^{2} − 4}

IV. Find J′(w) for

J(w) = \frac{w^{4} − 7w^{3} + 2w − 1}{w^{3/5}}